OpenStudy (anonymous):

Calculating the average rate of change for 1/(t-1) for t not equal to 1?

OpenStudy (anonymous):

Average rate of change? Over what interval?

OpenStudy (anonymous):

The first part of the problem is just to find the average rate of change. Fractions are breaking my brain. I know I have to put it in the formula f(x)=(x+h)-x/h

OpenStudy (anonymous):

Provided that the function is continuous over a given interval \([a,b]\), the average rate of change would be \[\frac{f(b)-f(a)}{b-a}\] or in your notation, we could let \(h=b-a~~\iff~~b=a+h\), so that you have \[\frac{f(a+h)-f(a)}{h}\]

OpenStudy (anonymous):

You need an interval to find an average. If we're doing it over \((-\infty, 1)\cup (1,\infty)\) that's tricky.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio You need an interval to find an average. If we're doing it over \((-\infty, 1)\cup (1,\infty)\) that's tricky. \(\color{blue}{\text{End of Quote}}\) Precisely

OpenStudy (anonymous):

I see what I did wrong...it's late, but thank you all :)